OpenStudy (anonymous):

What is the value of a(1) for a geometric sequence with a(4) = 40 and a(6) = 160? A. 2 B. 5 C. 8 D. 10

OpenStudy (campbell_st):

well you need to find the common ratio 1st \[160 = a_{1} \times r^{6 - 1}\] and the 2nd equation is \[40 = a_{1} \times r^{4 - 1}\] you may be able to guess the common ratio.... but rewriting the 1st equation \[160 = a_{1} \times r^3 \times r^2\] substituting the 2nd equation you get \[160 = 40 \times r^2\] solve for r... then substitute it into either equation to find the 1st term

OpenStudy (anonymous):

There is a way to work this out neatly, but I used to cheat on these questions and just input each answer into a calculator and fish for the correct answer. a_6=a_1*r^(n-1) 160 = a * r^(6-1) If you know r then you only need to solve for a we know here that 40 * r * r = 160 so we can solve for r... 40 * r * r /40 = 160/40 r * r = 4 r^2 = 4 SQRT(r^2) = SQRT(4) r = 2 so now we have 160 = a * 2^(5) and now can you solve for a?

OpenStudy (anonymous):

Thanks I got it!